OpenStudy (anonymous):

given f(x)+ (x)/(x+(2/X)) The derivative of the function is given by f'(x)= (Ax^2+Bx+c)/(x^2+D)^2 where A=__ B=__C=__D=__ Everytime I try to use the quotient rule I dont come up with the correct answer. How do I do this? Can you please show each step? Thank you

OpenStudy (anonymous):

typo the F(x)+ is supposed to be an = not a +

OpenStudy (helder_edwin):

\[ \large f(x)=\frac{x}{x+\frac{2}{x}} \] ??

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

then after some operations: \[ \large f(x)=\frac{x^2}{x^2+2} \] agree?

OpenStudy (anonymous):

how did you get that?

OpenStudy (helder_edwin):

add the fractions in the denominator. and then use this \[ \large \frac{a/b}{c/d}=\frac{ad}{bc} \]

OpenStudy (anonymous):

use quotient rule:\[f'(x)=\frac{1(x+\frac{2}{x})-x(1-\frac{2}{x^2})}{(x+\frac{2}{x})^2}\]go from there

OpenStudy (anonymous):

Yeah, better to simplify first as per @helder_edwin

OpenStudy (anonymous):

then use quotient rule:\[f'(x)=\frac{2x(x^2+2)-x^2(2x)}{(x^2+2)^2}\]go fro, there...

OpenStudy (anonymous):

thanks guys! I didnt simplify before I tried using the quotient rule so I think that's where i went wrong

OpenStudy (anonymous):

Yeah, if you don't simplify before hand it is easy to make a mistake...the calculation is much messier.